# Interpretable and Robust AI in EEG Systems: A Survey

### v3

@@ II-A3 Event-Related Desynchronization and Synchronization
Emotion Recognition Deap [11] EEG, EMG, EOG, BVP $32$ $19-37$ (Mean $26.9$)
@@
SEED-IV [13] EEG $15$ $20-24$
SEED-V [14] EEG, SMI $20$ N/A
Dreamer [15] EEG, ECG $23$ $20-33$
HCI-Tag [16] EEG, ECG, GSR, TEMP, RSP $30$ $19-40$ (Mean $26.1$)
@@
Imagined Emotion [20] EEG $31$ $18-38$
Fatigue Detection MEDT [21] EEG $27$ $22-28$
@@
Freiburg Seizure [26] EEG $21$ Adults and Children
@@
Temple University [28] EEG $80$ $6-56$
NMT [29] EEG N/A $1-90$
Sleep Monitoring Sleep-EDF [30] EEG, EMG, EOG $197$ $18-90$
@@
HMC [34] EEG, EMG, EOG $105$ $20-80$
CAP Sleep [35] EEG, ECG, EOG $102$ $18-70$
Sleep Cassette [36] EEG, EOG, EMG $100$ $18-65$
SIESTA [37] EEG, EOG, EMG, ECG $72$ $18-80$
Motor Imagery BCI IV 2a* [38] EEG, EOG $9$ N/A
BCI IV 2b* [38] EEG, EOG $9$ N/A
OpenMBI [39] EEG $54$ $24-45$
Stroke [40, 41] EEG $21$ Mean $54.2$
MOABB [42] EEG $104$ $18-55$
@@ III-A1 Layer-wise Relevance Propagation (LRP)
where $f_{j}^{(l-1)}$ represents the activation value of neuron $j$ in layer $(l-1)$, $w_{ji}^{(l)}$ denotes the weight between neuron $i$ in layer $l$ and neuron $j$ in layer $(l-1)$, and $z^{(l)}_{i}=\sum_{j}f_{j}^{(l-1)}w_{ji}^{(l)}$ denotes the sum of the inputs to neuron $i$ in layer $l$.
@@ III-C2 Fuzzy Inference System (FIS)
$R_{k}:\text{ IF }x_{1}\text{ is }A_{k1}\text{ AND }\cdots\text{ AND }x_{n}\text{ is }A_{kn}\text{ THEN }y\text{ is }B_{k}\text{, }$
@@ IV-B Human Variability
$D(\mathcal{S},\mathcal{T})=\sum_{i=1}^{n}||\mathbf{w}_{i}^{\mathcal{S}}-\mathbf{w}_{i}^{\mathcal{T}}||_{2}^{2},$ (21)
@@
$\min_{\mathbf{L},\mathbf{E}}||\mathbf{L}||_{*}+\lambda||\mathbf{E}||_{1}\quad\text{s.t.}\quad\mathbf{X}=\mathbf{L}+\mathbf{E},$ (22)

### v4

@@ II-A3 Event-Related Desynchronization and Synchronization
Emotion Recognition Deap [11] EEG, EMG, EOG, BVP $32$ 193719-37 (Mean $26.9$)
@@
SEED-IV [13] EEG $15$ 202420-24
SEED-V [14] EEG, SMI 2020 N/A
Dreamer [15] EEG, ECG $23$ 203320-33
HCI-Tag [16] EEG, ECG, GSR, TEMP, RSP $30$ 194019-40 (Mean $26.1$)
@@
Imagined Emotion [20] EEG $31$ 183818-38
Fatigue Detection MEDT [21] EEG $27$ 222822-28
@@
Freiburg Seizure [26] EEG 2121 Adults and Children
@@
Temple University [28] EEG $80$ 6566-56
NMT [29] EEG N/A 1901-90
Sleep Monitoring Sleep-EDF [30] EEG, EMG, EOG $197$ 189018-90
@@
HMC [34] EEG, EMG, EOG $105$ 208020-80
CAP Sleep [35] EEG, ECG, EOG $102$ 187018-70
Sleep Cassette [36] EEG, EOG, EMG $100$ 186518-65
SIESTA [37] EEG, EOG, EMG, ECG $72$ 188018-80
Motor Imagery BCI IV 2a* [38] EEG, EOG 99 N/A
BCI IV 2b* [38] EEG, EOG 99 N/A
OpenMBI [39] EEG $54$ 244524-45
Stroke [40, 41] EEG 2121 Mean $54.2$
MOABB [42] EEG $104$ 185518-55
@@ III-A1 Layer-wise Relevance Propagation (LRP)
where $f_{j}^{(l-1)}$ represents the activation value of neuron $j$ in layer (l1)(l-1), $w_{ji}^{(l)}$ denotes the weight between neuron $i$ in layer $l$ and neuron $j$ in layer (l1)(l-1), and $z^{(l)}_{i}=\sum_{j}f_{j}^{(l-1)}w_{ji}^{(l)}$ denotes the sum of the inputs to neuron $i$ in layer $l$.
@@ III-C2 Fuzzy Inference System (FIS)
[…]
@@ IV-B Human Variability
[…]
@@
[…]